3. A projectile leaves the ground with initial velocity of 30 m/s, at an angle of θ above the

horizontal. At maximum height, the velocity of the projectile is half of its initial velocity.

a) Calculate the angle, θ of the projectile motion [2]


b) Calculate the time, ttotal needed for the projectile to reach the ground. [3]


1
Expert's answer
2021-11-11T11:59:16-0500

Solution.

"v_0=30 m\/s;"

"v=0.5v_0;"

"\theta -?; t -?;"

"a) v_x=v_{0x}=v_0cos\theta\implies cos\theta=\dfrac{v_x}{v_{0x}}" ;

"cos\theta=\dfrac{0.5\sdot30 m/s}{30 m/s}=0.5; \theta=60^o;"

"b) t=\dfrac{2v_0sin\theta}{g};"

"t=\dfrac{2\sdot30 m/s\sdot0.866}{9.8 m/s^2}=5.3s;"

Answer: "a)\theta=60^o;"

"b)t=5.3s."